OpenStudy (anonymous):

**MEDAL + FAN** This is my last flashcard! :) One-eighth of a bookstore's magazines are sold on a Friday. If ¼ of the remaining magazines are sold the next day, what fractional part of the magazines remains at the end of the second day? A. 1/32 B. 1/8 C. 7/32 D. 21/32 <--- Formula? @gorv @Asevilla5

OpenStudy (gorv):

first day sold 1/8 remaining 1-1/8=7/8

OpenStudy (gorv):

sold on second = 7/8 * 1/4= 7/32 remain=7/8-7/32=21/32